how many liters of a 75% solution should be mixed with 55% solution to get 70 liters of a 63% solution?
Answer by checkley77(12844)

.75X+.55(70-X)=.63*70
.75X+38.5-.55X=44.1
.20X=44.1-38.5
.20X=5.6
X=5.6/.20
X=28 LITERS OF 75% SOLUTION.
70-28=42 LITERS OF 55% SOLUTION.
PROOF:
.75*28+.55*42=.63*70
21+23.1=44.1
44.1=44.1